# Crowdsourcing Ubiquitous Indoor Localization with Non-Cooperative Wi-Fi Ranging

### v1

@@ Abstract.
Indoor localization opens the path to potentially transformative applications. Although many indoor localization methods have been proposed over the years, they remain too impractical for widespread deployment in the real world. In this paper, we introduce PeepLoc, a deployable and scalable Wi-Fi-based solution for indoor localization that relies only on pre-existing devices and infrastructure. Specifically, PeepLoc works on any mobile device with an unmodified Wi-Fi transceiver and in any indoor environment with a sufficient number of Wi-Fi access points (APs) and pedestrian traffic. At the core of PeepLoc is (a) a mechanism which allows any Wi-Fi device to obtain non-cooperative time-of-flight (ToF) to any Wi-Fi AP and (b) a novel bootstrapping mechanism that relies on pedestrian dead reckoning (PDR) and crowdsourcing to opportunistically initialize pre-existing APs as anchor points within an environment. We implement PeepLoc using commodity hardware and evaluate it extensively across 4 campus buildings. We show PeepLoc leads to a mean and median positional error of 3.41 m and 3.06 m respectively, which is superior to existing deployed indoor localization systems and is competitive with commodity GPS in outdoor environments.
@@ 1. Introduction
Figure 1. Non-cooperative time-of-flight measurements (i.e. one-way ranging) enables ranging between any Wi-Fi AP and client pair without specialized hardware requirements.
@@
(1) $t_{\text{RTT}}=\frac{2d}{c}+t_{\text{proc}}$
@@
In this paper, we propose PeepLoc, which addresses these challenges through a combination of new techniques and system design. At the core of PeepLoc is (a) non-cooperative Wi-Fi ranging (abedi_non-cooperative_2022, ), which enables any Wi-Fi enabled device to conduct ranging measurements to any other Wi-Fi enabled device , and (b) accurate pedestrian dead-reckoning capabilities on modern smartphones, which we use to be able to harvest valuable information from large numbers of smartphone users regularly entering indoor environments. PeepLoc leverages these to bring the crowdsourcing framework to Wi-Fi ranging-based indoor positioning systems, enhancing accuracy, adaptability, and scalability. In the following, we elaborate on three key aspects of PeepLoc.
  • Crowdsourcing APs with Non-Cooperative Wi-Fi Ranging. To overcome the scalability limitations of site surveying, the innaccuracies of RSS-based methods, and the hardware assumptions of GPS-based AP localization (e.g., OpenLocate), PeepLoc opportunistically collects non-cooperative Wi-Fi ranging measurements from smartphones on pedestrians as they enter environments in order to locate APs. We use non-cooperative Wi-Fi ranging over FTM as it works with any 802.11-compliant device-responder pair, meaning PeepLoc can potentially use any AP in the environment. The core idea is as follows. While a pedestrian is outdoors, PeepLoc can geolocate the smartphone using GPS, however upon entering a building, GPS is lost. PeepLoc detects the point at which GPS is lost and then switches to PDR (pedestrian dead reckoning) to continue getting accurate positioning (up to a certain point where intolerable drift will accumulate). While the trajectory of the smartphone is known, PeepLoc uses non-cooperative Wi-Fi ranging measurements to multilaterate and thus geolocate the positions of APs encountered. Over time, repeated visits by multiple users from diverse paths improve localization accuracy via spatial averaging. This approach avoids the need for dedicated setup efforts, leverages everyday motion for mapping, and naturally captures environmental diversity, improving AP localization in cluttered or dynamic environments.
@@
  • We contribute a toolkit for collecting non-cooperative Wi-Fi ranging measurements using commodity ESP32s. These are available at the following  [^1] https://github.com/ConnectedSystemsLab/esp32_wipeep_ros. We hope this drives further research into Wi-Fi ranging-based indoor positioning systems.
2. Background
@@
(b) One-way Wi-Fi ranging.
Figure 4. Comparison between 802.11mc (FTM) protocol and one-way Wi-Fi ranging. Note that the FTM protocol requires cooperation of both devices, whereas one-way ranging works with any standards-compliant Wi-Fi device.
(a) RTT vs distance in one-way Wi-Fi ranging.
@@
Wi-Fi Ranging. One of the primary methods for obtaining ranging estimates from Wi-Fi is the Fine Time Measurement (FTM) protocol, standardized in IEEE 802.11mc. FTM is a two-way ranging process that requires cooperation between an initiating device (typically a client station) and a responding device (usually an access point) (see Fig. 4(a)). The process begins when the initiator sends an FTM Request to the responder. Upon agreement, the responder transmits one or more FTM frames, each containing precise timestamps indicating when the frame was sent and when it was received. The initiator responds to each FTM frame with an acknowledgment (ACK) and records the corresponding timestamps. By analyzing these timestamps, the initiator calculates the round-trip time (RTT) of the signals, from which it derives the time-of-flight (ToF) and thus estimates the distance to the responder. While FTM provides accurate ranging capabilities, its adoption is limited, as it requires both devices to support the 802.11mc standard. As of now, only certain manufacturers, such as Aruba and Google, offer devices with 802.11mc support, making widespread deployment of FTM-based ranging systems less common.
Alternatively, non-cooperative one-way ranging—also known as polite Wi-Fi—offers a lightweight method that works with any standards-compliant Wi-Fi device, without requiring support for 802.11mc (see Fig. 4(b)). In this approach, device A transmits a data packet to device B and records the round-trip time (RTT) upon receiving an acknowledgment (ACK). To extract the time-of-flight (ToF), device A must subtract the processing delay incurred at device B. However, since device B does not explicitly report its processing time, estimating this delay is the central challenge of one-way ranging. WiPeep (abedi_non-cooperative_2022, ) addresses this challenge by collecting one-way RTT measurements from multiple spatially distributed vantage points. By exploiting spatial diversity and the broadcast nature of Wi-Fi, it infers the hidden processing delay and thereby recovers the ToF, enabling accurate distance estimation without device cooperation.
@@ 3.2. Design Choices
Advantages of Wi-Fi Ranging. For both AP and client localization, our system uses Wi-Fi ranging measurements over alternatives such as RSS and AoA for the following reasons. First, RSS ranging measurements are inaccurate due to coarse spatial resolution and sensitivity to noise/interference. Second, AoA-based triangulation would require multiple antennas with known geometries at the client (necessitating hardware overhead) and would rely on accurate device orientation to pinpoint the location of APs (limiting overall accuracy). By leveraging Wi-Fi ToF measurements, we can sidestep the shortcomings of RSS-based methods and the excessive requirements of AoA-based methods.
@@ 4.2. AP Discovery & Localization
(2) $t_{i}=\frac{2d_{i}}{c}+c_{j}+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}),$
@@
(3) $\ell(x,y,c)=\sum_{i=1}^{n}\left\|t_{i}-\left(\frac{2}{c}\|(x_{i},y_{i})-(x,y)\|+c\right)\right\|^{2}.$
@@
(4) $(\hat{x}_{j},\hat{y}_{j},\hat{c}_{j})=\underset{(x,y,c)}{\arg\min}\sum_{i=1}^{n}\left\|t_{i}-\left(\frac{2}{c}\|(x_{i},y_{i})-(x,y)\|+c\right)\right\|^{2}$
@@ 4.3. Client Localization
(5) $t_{\text{RTT}}=\alpha\cdot d+t_{\text{proc}},$
@@
(6) $(\hat{x},\hat{y})=\underset{(x,y)}{\arg\min}\sum_{i=1}^{n}\left\|t_{i}-\left(\alpha_{b_{i}}\cdot\|(x,y)-(x_{b_{i}},y_{b_{i}})\|+c_{b_{i}}\right)\right\|^{2}.$
@@ 6.1. Methodology
We evaluate PeepLoc across 5 different floorplan configurations spanning 4 different campus buildings (A,B,C,D), summarize in Table. 1. Each floorplan configuration consists of a physical building floor and as well as an arrangement of APs within it (we use a combination of commodity 802.11ac APs manufactured by Netgear R6300v2 and ASUS RT-AC86U). For simplicity, we set all APs in the environment to the same channel as the ESP32s. We carry out data collection inside and around the campus buildings. During data collection, we collect several traces by moving the sensors around while they are mounted on a pedestrian. This effectively simulates the trajectories of pedestrians entering the building. Furthermore, we also sample a set of test points in the building to assess location accuracy. For each test point, we collect around a few hundred Wi-Fi readings from the tandem ESP32s. We discard test points in which we detect less than 3 APs (in which case, the multi-lateration problem is under-determined). To compare PeepLoc’s performance to state-of-the-art baselines from both industry and academia, we also collect data for said baselines at the same test points.
Table 1. Building floorplan dimensions and data collection summary
A $94.17$ $37.40$ $10$ $11$ $5$ $4$
B $37.69$ $74.78$ $9$ $10$ $8$ 00
C $26.12$ $50.49$ $9$ $6$ $8$ 00
D1 $30.24$ $78.45$ $11$ $10$ $9$ 00
D2 $30.24$ $78.45$ $14$ $15$ $4$ $3$
@@ 6.2. AP & Client Location Accuracy
6.3. Ablation Study
We conduct an ablation study on the per-AP ranging model by comparing the localization accuracy with and without the per-AP ranging model enabled (in which case we use a fixed slope of $\frac{c}{2}$). We summarize the results in Fig. 9(c). As the graph shows, the per-AP ranging model is crucial for accurate location estimation.
@@ 6.3. Ablation Study
6.4. Comparison with Baselines
Finally, we compare PeepLoc with currently deployed indoor localization systems as well as state-of-the-art fingerprinting baselines.
  • CoreLocation. This is the geolocation service on iOS. This service uses Wi-Fi, GPS, Bluetooth, magnetometer, barometer, cellular hardware, and ranging measurements to nearby iBeacons.
  • Fused Location Provider. This is the Android geolocation service that combines various location sources, such as GPS, Wi-Fi, cellular networks, and ranging measurements towards nearby 802.11mc FTM-capable devices (e.g. Google Nest, Aruba APs). We note that all of the campus buildings that we run experiments in (i.e. all floorplans in Table. 1) have an enterprise-grade deployment of FTM-capable Aruba AP-635 (hpe_ap635_2025, ) running OpenLocate (troymart_its_2023, ).
  • DumbLoc (dumbloc_sensors2024, ). A state-of-the-art machine learning-based Wi-Fi fingerprinting framework using XGBoost and Random Forest models for indoor localization. It achieves meter-level accuracy through RSSI pattern matching without requiring parameter optimization or pre-mapped AP locations, designed for GPS-denied environments with existing Wi-Fi infrastructure. We configure DumbLoc to use the 5 APs with strongest signal strength for feature extraction.

### v2

@@ Abstract.
Indoor localization opens the path to potentially transformative applications. Although many indoor localization methods have been proposed over the years, they remain too impractical for widespread deployment in the real world. In this paper, we introduce PeepLoc, a deployable and scalable Wi-Fi-based solution for indoor localization that relies only on existing devices and infrastructure. Specifically, PeepLoc works on any mobile device with an unmodified Wi-Fi transceiver and in any indoor environment with a sufficient number of Wi-Fi access points (APs) and pedestrian traffic. At the core of PeepLoc is (a) a mechanism which allows any Wi-Fi device to obtain non-cooperative time-of-flight (ToF) to any Wi-Fi AP and (b) a novel bootstrapping mechanism that relies on pedestrian dead reckoning (PDR) and crowdsourcing to opportunistically initialize pre-existing APs as anchor points within an environment. We implement PeepLoc using commodity hardware and evaluate it extensively across 4 campus buildings. We show PeepLoc leads to a mean and median positional error of 3.41 m and 3.06 m respectively, which is superior to existing state-of-the-art deployed indoor localization systems and is competitive with commodity GPS in outdoor environments.
@@ 1. Introduction
Figure 1. Non-cooperative time-of-flight measurements (i.e. one-way ranging) enables ranging between any Wi-Fi AP and client pair without specialized hardware requirements or modifications.
@@
(1) $t_{\text{RTT}}=\frac{2d}{c}+t_{\text{proc}}\vspace{0.1in}$
@@
In this paper, we propose PeepLoc, which addresses these challenges through a combination of new techniques and system design. At the core of PeepLoc is (a) non-cooperative Wi-Fi ranging (abedi_non-cooperative_2022, ), which enables any Wi-Fi enabled device to conduct ranging measurements to any other Wi-Fi enabled device, and (b) accurate pedestrian dead-reckoning capabilities on modern smartphones, which we use to be able to harvest valuable information from large numbers of smartphone users regularly entering indoor environments. PeepLoc leverages these to bring the crowdsourcing framework to Wi-Fi ranging-based indoor positioning systems, enhancing accuracy, adaptability, and scalability. Moreover, while some prior indoor localization systems report high accuracy in controlled settings, they typically rely on specialized hardware, custom firmware, or labor-intensive site survey that severely limit their scalability and practicality in real-world environments. In contrast, PeepLoc is designed to operate entirely on commodity smartphones and unmodified Wi-Fi infrastructure. It does not require PHY-layer access, protocol support like 802.11mc/az, or prior knowledge of infrastructure layout. This “zero-assumption” design makes PeepLoc uniquely deployable at scale, turning ubiquitous Wi-Fi deployments and everyday pedestrian motion into a crowdsourced indoor localization system. In the following, we elaborate on three key aspects of PeepLoc.
  • Crowdsourcing APs with Non-Cooperative Ranging. To overcome the scalability limitations of site surveying, the innaccuracies of RSS-based methods, and the hardware assumptions of GPS-based AP localization (e.g., OpenLocate), PeepLoc opportunistically collects non-cooperative Wi-Fi ranging measurements from smartphones on pedestrians as they enter environments in order to locate APs. We use non-cooperative Wi-Fi ranging over FTM as it works with any 802.11-compliant device-responder pair, meaning PeepLoc can potentially use any AP in the environment. The core idea is as follows. While a pedestrian is outdoors, PeepLoc can geolocate the smartphone using GPS, however upon entering a building, GPS is lost. PeepLoc detects the point at which GPS is lost and then switches to PDR (pedestrian dead reckoning) to continue getting accurate positioning (up to a certain point where intolerable drift will accumulate). While the trajectory of the smartphone is known, PeepLoc uses non-cooperative Wi-Fi ranging measurements to multilaterate and thus geolocate the positions of APs encountered. Over time, repeated visits by multiple users from diverse paths improve localization accuracy via spatial averaging. This approach avoids the need for dedicated setup efforts, leverages everyday motion for mapping, and naturally captures environmental diversity, improving AP localization in cluttered or dynamic environments.
@@
Table 1. Comparison of PeepLoc against prior work/existing systems.
System No AP Cooperation Needed No PHY Access Required No Prior AP Locations or Map Needed No Special Hardware Needed Data Collection Overhead Deployability
RADAR/Horus/Zee (bahl2000radar, ; youssef2005horus, ; chintalapudi_zee_2012, ) High Medium
ArrayTrack/SpotFi (xiong_arraytrack_2013, ; kotaru2015spotfi, ) Medium Low
Chronos (vasisht2016chronos, ) Low Low
Horn’s System (horn_indoor_2022, ) Medium Low
Wi-Fi RTT (802.11mc/az) (IEEE802.11mc2016, ; IEEE802.11az2022, ; ibrahim2018rtt, ) Low Low
PeepLoc (Ours) Low High
@@
  • We contribute a toolkit for collecting non-cooperative Wi-Fi ranging measurements on commodity ESP32s. These are available at  [^1] https://github.com/ConnectedSystemsLab/esp32_wipeep_ros. We hope this drives further research into Wi-Fi ranging-based indoor positioning systems.
2. Background & Related Work
2.1. Background
@@
(b) Non-Cooperative Wi-Fi ranging.
Figure 4. Comparison between 802.11mc (FTM) protocol and one-way Wi-Fi ranging. Note that the FTM protocol requires cooperation of both devices, whereas non-cooperative ranging works with any standards-compliant Wi-Fi device.
(a) RTT vs distance in non-cooperative Wi-Fi ranging.
@@
Wi-Fi Ranging. Wi-Fi-based ranging primarily relies on the Fine Time Measurement (FTM) protocol, standardized in IEEE 802.11mc (IEEE802.11mc2016, ) and extended in the newer IEEE 802.11az (IEEE802.11az2022, ). These protocols support both two-way and one-way RTT time-of-flight (ToF) estimation by exchanging timestamped packets between an initiating device (typically a client station like a smartphone) and a responding device (usually a Wi-Fi access point) (see Fig. 4(a)). In the standard FTM procedure, the initiator sends a request package at timestamp $\textbf{t}_{1}$, the responder marks the reception timestamp and replies an ACK package containing precise reception timestamp $\textbf{t}_{2}$ and transmission timestamp $\textbf{t}_{3}$. Upon receiving the ACK package from the responsder at $\textbf{t}_{4}$, the precise round-trip time can be estimated by Eq. 2. By analyzing the round-trip time (RTT) across these frames and acknowledgments (ACKs), the initiator can estimate ToF and thus compute distance.
(2) $\text{ToF}_{\text{2-sided}}={(t_{4}-t_{1})-(t_{3}-t_{2})}$
While FTM-based ranging offers high accuracy, its adoption remains limited due to its dependence on hardware and firmware support at both ends of the link. Only a subset of commercial devices—primarily high-end smartphones (e.g., Google Pixel, Samsung Galaxy S series) and APs from vendors like Aruba—implement the required standards. As a result, widespread deployment is constrained by specialized hardware requirements. For responder like legacy Wi-Fi access point which lacks the precise timing capability, the FTM protocol degrades to one-way RTT, where responder doesn’t provide the reception and transmission timestamps. In such cases, the initiator can only approximate the round-trip time as Eq. 3 unless the AP’s turn-around offset can be estimated.
(3) $\text{ToF}_{\text{1-sided}}={(t_{4}-t_{1})}$
In contrast, non-cooperative ranging presents a lightweight alternative that operates on any standards-compliant Wi-Fi device without relying on 802.11mc/az (see Fig. 4(b)). In this mode, device A transmits a dummy data packet (Null Data Packet) to device B and measures the RTT based on the ACK response from device B. However, since the receiver’s internal processing delay is not reported, estimating ToF becomes challenging. Systems like WiPeep (abedi_non-cooperative_2022, ) address this by leveraging spatial diversity and passive RTT measurements from multiple vantage points to infer hidden delays and accurately estimate distance—enabling ranging without protocol support or device cooperation.
@@
2.2. Related Work
Localization with Wi-Fi Fingerprinting. Fingerprinting remains one of the most widely adopted techniques for indoor localization. It leverages ambient signal measurements from various sensors (e.g. Wi-Fi, BLE, magnetometer) to create a map of the environment. The seminal RADAR system (bahl2000radar, ) demonstrated that Wi-Fi RSS fingerprints could localize users with median errors of 2–3 meters in an office environment without additional hardware infrastructure. Subsequent work has improved on this framework using better statistical models e.g. Horus (youssef2005horus, ), including more sensors (wang_no_2012, ), incorporating crowdsourcing e.g. Zee (chintalapudi_zee_2012, ), and incorporating deep learning (nowicki2017wifi, ). Fingerprinting has resulted in many commercial-grade crowdsourced systems (e.g. Apple’s indoor maps program (apple_indoor_maps, ) and Tencent’s indoor localization system (ni_experience_2022, )). Despite their popularity, fingerprinting systems have well known challenges. First, they require costly and time-consuming data collection. Second, they require that floorplans or maps of the environment are known beforehand. Finally, their performance degrades in dynamic environments.
Localization With Wi-Fi PHY Layer Information. Another well-researched class of localization systems infer location from observing the physical channel at either the client device or the anchor points (Wi-Fi APs). Representative examples include ArrayTrack (xiong_arraytrack_2013, ), SpotFi (kotaru2015spotfi, ), and and Chronos (vasisht2016chronos, ). ArrayTrack (xiong_arraytrack_2013, ) uses PHY information from antenna arrays at MIMO APs to measure AoA towards clients and triangulate their location. SpotFi (kotaru2015spotfi, ) extends this framework by adding MUSIC super-resolution and clustering to find the LOS path. Breaking away from the AoA framework, Chronos (vasisht2016chronos, ) instead measures channel estimates across multiple frequency bands to estimate ToF between a device and AP pair. Although these methods lead to competitive localization accuracy, they suffer from several drawbacks. First, they require access to low-level PHY data and customized firmware. Second, they require known AP locations and orientations in the case of AoA-based systems. These factors limit their practical deployment.
Localization with Cooperative Ranging. To overcome the disadvantages of fingerprinting and PHY-reliant methods, time-of-flight (ToF) based methods have been proposed beginning with the first two-way Wi-Fi ranging standard IEEE 802.11mc (and later 802.11az). These protocols require implementation across both Wi-Fi APs and clients to enable out-of-the-box estimation of round-trip propagation delay. Ibrahim et al. evaluated the feasibility of Wi-Fi RTT on commercial smartphones and achieve sub-2 meter accuracy in line-of-sight indoor settings (ibrahim2018rtt, ). Although two-way Wi-Fi ranging has high accuracy, it is supported only by a few devices and AP manufacturers on the market (e.g. Aruba Open Locate (troymart_its_2023, )), limiting its widespread deployment.
Localization with Non-Cooperative Wi-Fi Ranging. Unlike two-way (cooperative) ranging methods, non-cooperative ranging methods do not rely on cooperation at the AP. These include our system, Wi-Peep (abedi_non-cooperative_2022, ), and Horn’s system (horn_indoor_2022, ). Wi-Peep tackles the problem of localizing devices inside an indoor environment using a GPS-equipped drone flying around the building. Since the drone functions as the anchor point for localization in this case, it is not suitable as a permanently deployed localization system. The system proposed by Horn (horn_indoor_2022, ) uses non-cooperative Wi-Fi ToF as an indoor localization system, but assumes the locations of the APs within the environment are known beforehand. By contrast, our system is capable of automatically locating where the APs are inside an environment through crowdsourcing.
PeepLoc in Context. As summarized in Table 1, PeepLoc stands out from prior work by eliminating common deployment barriers across both cooperative and non-cooperative localization systems. Unlike fingerprinting approaches (RADAR(bahl2000radar, ), Horus(youssef2005horus, ), Zee (chintalapudi_zee_2012, )) that require labor-intensive data collection and known maps, or PHY-dependent systems (ArrayTrack (xiong_arraytrack_2013, ), SpotFi (kotaru2015spotfi, ), Chronos (vasisht2016chronos, )) that rely on firmware or hardware modifications, PeepLoc operates without any specialized support from the access points, physical layer access, or knowledge of infrastructure layout. Moreover, unlike prior one-way ranging systems such as WiPeep (abedi_non-cooperative_2022, ) or Horn’s system (horn_indoor_2022, ), PeepLoc does not rely on external infrastructure (e.g., drones or GPS) or assume pre-calibrated AP locations. Instead, it introduces a novel system design that is fully software-defined and self-contained: all ranging and inference are performed using standard Wi-Fi frames and a commodity smartphone alone. This softwarized approach, requiring no firmware changes or hardware capabilities beyond what is already present in 802.11-compliant devices, enables true plug-and-play deployability. In doing so, PeepLoc advances the state of the art by bridging the gap between deployability, compatibility, and accurate ranging—all within an unmodified wireless stack.
@@ 3.2. Design Choices
Advantages of Wi-Fi Ranging. For both AP and client localization, our system uses Wi-Fi ranging measurements over alternatives such as RSS and AoA for the following reasons. First, RSS ranging measurements are inaccurate due to coarse spatial resolution and sensitivity to noise/interference/Wi-Fi chip. Second, AoA-based triangulation would require multiple antennas with known geometries at the client (necessitating hardware overhead) and would rely on accurate device orientation to pinpoint the location of APs (limiting overall accuracy). By leveraging Wi-Fi ToF measurements, we can sidestep the shortcomings of RSS-based methods and the excessive requirements of AoA-based methods.
@@ 4.2. AP Discovery & Localization
(4) $t_{i}=\frac{2d_{i}}{c}+c_{j}+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}),\vspace{0.1in}$
@@
(5) $\ell(x,y,c)=\sum_{i=1}^{n}\left\|t_{i}-\left(\frac{2}{c}\|(x_{i},y_{i})-(x,y)\|+c\right)\right\|^{2}.\vspace{0.1in}$
@@
(6) $\underset{(x,y,c)}{\arg\min}\sum_{i=1}^{n}\left\|t_{i}-\left(\frac{2}{c}\|(x_{i},y_{i})-(x,y)\|+c\right)\right\|^{2}\\ \cdot\mathds{1}\{b_{i}=b_{j}\}.\vspace{0.1in}$
@@ 4.3. Client Localization
(7) $t_{\text{RTT}}=\alpha\cdot d+t_{\text{proc}},\vspace{0.1in}$
@@
(8) $(\hat{x},\hat{y})=\underset{(x,y)}{\arg\min}\sum_{i=1}^{n}\left\|t_{i}-\left(\alpha_{b_{i}}\cdot\|(x,y)-(x_{b_{i}},y_{b_{i}})\|+c_{b_{i}}\right)\right\|^{2}.\vspace{0.1in}$
@@ 5. Implementation
Table 2. Building floorplan dimensions and data collection summary
A $94.17$ $37.40$ $10$ $11$ $5$ $4$
B $37.69$ $74.78$ $9$ $10$ $8$ 00
C $26.12$ $50.49$ $9$ $6$ $8$ 00
D1 $30.24$ $78.45$ $11$ $10$ $9$ 00
D2 $30.24$ $78.45$ $14$ $15$ $4$ $3$
@@ 6.1. Methodology
We evaluate PeepLoc across 5 different floorplan configurations spanning 4 different campus buildings (A,B,C,D), summarize in Table. 2. Each floorplan configuration consists of a physical building floor and as well as an arrangement of APs within it (we use a combination of commodity 802.11ac APs manufactured by Netgear R6300v2 (netgear:R6300v2, ) and ASUS RT-AC86U (asus:RT-AC86U, )). For simplicity, we set all APs in the environment to the same channel as the ESP32s. We carry out data collection inside and around the campus buildings. During data collection, we collect several traces by moving the sensors around while they are mounted on a pedestrian. This effectively simulates the trajectories of pedestrians entering the building. Furthermore, we also sample a set of test points in the building to assess location accuracy. For each test point, we collect around a few hundred Wi-Fi readings from the tandem ESP32s. We discard test points in which we detect less than 3 APs (in which case, the multi-lateration problem is under-determined). To compare PeepLoc’s performance to state-of-the-art baselines from both industry and academia, we also collect data for said baselines at the same test points.
@@ 6.2. AP & Client Location Accuracy
6.3. Comparison with State-of-the-Art
We compare PeepLoc with currently deployed indoor localization systems as well as state-of-the-art fingerprinting baselines.
  • CoreLocation (apple:corelocation, ). This is the geolocation service on iOS. This service uses Wi-Fi, GPS, Bluetooth, magnetometer, barometer, cellular hardware, and Bluetooth-based ranging measurements to nearby iBeacons.
  • Fused Location Provider (google:fusedlocationprovider, ). This is the Android geolocation service that combines various location sources, such as GPS, Wi-Fi, cellular networks, and ranging measurements towards nearby 802.11mc FTM-capable devices (e.g. Google Nest, Aruba APs). We note that all of the campus buildings that we run experiments in (i.e. all floorplans in Table. 2) have an enterprise-grade deployment of FTM-capable Aruba AP-635 (hpe_ap635_2025, ) running OpenLocate (troymart_its_2023, ).
  • DumbLoc (dumbloc_sensors2024, ). A state-of-the-art machine learning-based Wi-Fi fingerprinting framework using XGBoost (chen2016xgboost, ) and Random Forest models for indoor localization. It achieves meter-level accuracy through RSSI pattern matching without requiring parameter optimization or pre-mapped AP locations, designed for GPS-denied environments with existing Wi-Fi infrastructure. We configure DumbLoc to use the 5 APs with strongest signal strength for feature extraction.